# Large Language Models are Good Prompt Learners for Low-Shot Image Classification

### v1

@@ Abstract
Low-shot image classification, where training images are limited or inaccessible, has benefited from recent progress on pre-trained vision-language (VL) models with strong generalizability, e.g. CLIP. Prompt learning methods built with VL models generate text features from the class names that only have confined class-specific information. Large Language Models (LLMs), with their vast encyclopedic knowledge, emerge as the complement. Thus, in this paper, we discuss the integration of LLMs to enhance pre-trained VL models, specifically on low-shot classification. However, the domain gap between language and vision blocks the direct application of LLMs. Thus, we propose LLaMP, Large Language Models as Prompt learners, that produces adaptive prompts for the CLIP text encoder, establishing it as the connecting bridge. Experiments show that, compared with other state-of-the-art prompt learning methods, LLaMP yields better performance on both zero-shot generalization and few-shot image classification, over a spectrum of 11 datasets.
@@ 2 Related Work
Zero-Shot Learning (ZSL). ZSL stands in contrast to traditional fully-supervised paradigms. Instead of relying on direct visual training samples, it leverages side information that can be drawn from a multitude of non-visual domains, including attributes [22], word embeddings [40, 36], and descriptive texts [34]. Zhang et al. [46] designed an embedding model to bridge the gap between seen and unseen categories. Concurrently, studies like [4, 44, 50] have spotlighted that generative models can produce features for unseen categories. Moreover, Graph Convolution Networks (GCN) [20] has been explored in research such as [40, 16] to for further generalization.
@@ 3.2 Adaptive Prompt Learning with LLMs
The goal of prompt tuning is to find an set of optimal prompts $\bm{p}=\{\bm{p}_{v},\bm{p}_{t}\}$ which maximizes the log likelihood of $P(x,y|\theta_{\mathcal{F}},\theta_{\mathcal{G}})$ over target downstream distribution $(\bm{x},\bm{y})\sim(\bm{X},\bm{Y})$:
@@
To overcome these problems, we propose to learn a meta function on the language side $\bm{p}_{t}=\Theta(y)$ which can adaptively estimate the optimal prompt for each category. An intuitive way to estimate proper prompts $\bm{p}$ for category name $y$ is to take advantage of the knowledge of the pre-trained Large Lanuage Models (LLM) $\mathcal{D}$ and extract discriminative descriptions of category $y$. For example, given the input text $\bm{z}$:“Describe {y}”,
@@
while $p_{i}$ being sequencially generated by $\mathcal{D}$ such that
@@
where $\mathcal{D}^{(i)}$ is the $i$-th forward iteration of $\mathcal{D}$, and $\mathcal{M}$ maps continues hidden states into discrete language tokens. To accerelate the process and to obtain $p$ in one pass, we approximate the above process with $K$ learnable prompts $\bm{p}_{l}=\{\theta_{1},...,\theta_{K}\}$ so that
@@ 3.3 The LLaMP Framework
$\bm{\tilde{y}_{l}}^{i}=\{\bm{t}_{bos}^{i},\bm{p}^{i}_{t},\bm{t}_{1}^{i},\bm{t}_{2}^{i},\dots,\bm{t}_{L}^{i},\bm{t}_{eos}^{i},\bm{\tilde{h}_{l}^{i}}\}$ (8)
@@ 4 Experiments
Table 1: Comparison with state-of-the-art methods on base-to-novel generalization. LLaMP shows strong generalization results over previous approaches on 11 image classification tasks. Absolute gains over PSRC are indicated in blue. $*{}^{*}$ is trained with ViT-B/32 image encoder instead of ViT-B/16.
@@ 4.1 Experiment Setup
CLIP [32] 78.79 67.31 95.43 85.34 80.44 97.37 82.90 45.36 73.28 69.96 87.21 82.11
CoOp [49] 79.89 71.87 95.57 91.87 83.07 97.07 84.20 43.40 74.67 69.87 84.93 82.23
CoCoOp [48] 74.90 70.83 95.16 93.34 71.57 87.84 87.25 31.21 72.15 63.04 73.32 78.14
MaPLe [18] 81.79 72.33 96.00 92.83 83.57 97.00 85.33 48.40 75.53 71.33 92.33 85.03
PSRC [19] 82.87 73.17 96.07 93.67 83.83 97.60 87.50 50.83 77.23 72.73 92.43 86.47
LLaMP 83.81 73.49 97.08 94.21 86.07 98.06 87.62 56.07 77.02 74.17 91.31 86.84
Table 2: Few shot classification results with 16 shots.
@@ 4.3 Ablation Study
Noun phrases are parsed from the LLM’s response the prompt of “Describe [STH]”. We then use the template, “A photo of [STH] with [NP]” to generate the NP-augmented text embedding for CLIP. We take the average of all augmented embeddings for classification. In Tab. 3 we show that even ordinary CLIP can benefit from incorporating LLMs’ knowledge.
@@
four engines, turboprop aircraft, large vertical fin
@@
a cluster, rectangular structures, flat roofs, straight lines
@@
Visualizations. In Fig. 4, we visualize the gradient heatmap of input images from FGVCAircraft and EuroSAT, through GradCAM [35]. The figure shows that, with the knowledge from an LLM, LLaMP is able to capture distinctive features that matches LLM’s description.
@@ 5 Conclusion
Appendix Overview
Here, we provide extra discussions that did not fit in the main paper due to the space limitation, including i) Ablation study on the textual augmentation described in Sec 3.3; ii) Few-shot classification results on 8/4/2/1 shots with comparisons with previous methods.
@@
In addition to the 16-shot classification result reported in Tab. 2, we present few-shot classification results with with 8/4/2/1 numbers of shots in Tab. 9 and compare LLaMP against previous baseline models.

### v2

@@ Abstract
Low-shot image classification, where training images are limited or inaccessible, has benefited from recent progress on pre-trained vision-language (VL) models with strong generalizability, e.g. CLIP. Prompt learning methods built with VL models generate text features from the class names that only have confined class-specific information. Large Language Models (LLMs), with their vast encyclopedic knowledge, emerge as the complement. Thus, in this paper, we discuss the integration of LLMs to enhance pre-trained VL models, specifically on low-shot classification. However, the domain gap between language and vision blocks the direct application of LLMs. Thus, we propose LLaMP, Large Language Models as Prompt learners, that produces adaptive prompts for the CLIP text encoder, establishing it as the connecting bridge. Experiments show that, compared with other state-of-the-art prompt learning methods, LLaMP yields better performance on both zero-shot generalization and few-shot image classification, over a spectrum of 11 datasets. Code will be made available at: https://github.com/zhaohengz/LLaMP.
@@ 2 Related Work
Zero-Shot Learning (ZSL). ZSL stands in contrast to traditional fully-supervised paradigms. Instead of relying on direct visual training samples, it leverages side information that can be drawn from a multitude of non-visual domains, including attributes [22], word embeddings [40, 36], and descriptive texts [34]. Zhang et al. [46] designed an embedding model to bridge the gap between seen and unseen categories. Concurrently, studies like [4, 44, 50] have spotlighted that generative models can produce features for unseen categories. Moreover, Graph Convolution Networks (GCN) [20] has been explored in research such as [40, 16] for further generalization.
@@ 3.2 Adaptive Prompt Learning with LLMs
The goal of prompt tuning is to find a set of optimal prompts $\bm{p}=\{\bm{p}_{v},\bm{p}_{t}\}$ which maximizes the log likelihood of $P(x,y|\theta_{\mathcal{F}},\theta_{\mathcal{G}})$ over target downstream distribution $(\bm{x},\bm{y})\sim(\bm{X},\bm{Y})$:
@@
To overcome these problems, we propose to learn a meta function on the language side $\bm{p}_{t}=\Theta(y)$ which can adaptively estimate the optimal prompt for each category. An intuitive way to estimate proper prompts $\bm{p}$ for category name $y$ is to take advantage of the knowledge of the pre-trained Large Language Models (LLM) $\mathcal{D}$ and extract discriminative descriptions of category $y$. For example, given the input text $\bm{z}$:“Describe {y}”,
@@
while $p_{i}$ being sequentially generated by $\mathcal{D}$ such that
@@
where $\mathcal{D}^{(i)}$ is the $i$-th forward iteration of $\mathcal{D}$, and $\mathcal{M}$ maps continuous hidden states into discrete language tokens. To accelerate the process and to obtain $p$ in one pass, we approximate the above process with $K$ learnable prompts $\bm{p}_{l}=\{\theta_{1},...,\theta_{K}\}$ so that
@@ 3.3 The LLaMP Framework
$\bm{\tilde{y}_{l}}^{i}=\{\bm{t}_{bos}^{i},\bm{p}^{i}_{t},\bm{t}_{1}^{i},\bm{t}_{2}^{i},\dots,\bm{t}_{L}^{i},\bm{\tilde{h}_{l}^{i}},\bm{t}_{eos}^{i}\}$ (8)
@@ 4 Experiments
Table 1: Comparison with state-of-the-art methods on base-to-novel generalization. LLaMP shows strong generalization results over previous approaches on 11 image classification tasks. Absolute gains over PSRC are indicated in blue. KAPT is trained with ViT-B/32 image encoder instead of ViT-B/16.
@@ 4.1 Experiment Setup
CLIP [32] 78.79 (65.02) 67.31 95.43 85.34 80.44 97.37 82.90 45.36 73.28 69.96 87.21 82.11
CoOp [49] 79.89 (73.82) 71.87 95.57 91.87 83.07 97.07 84.20 43.40 74.67 69.87 84.93 82.23
CoCoOp [48] 74.90 (70.70) 70.83 95.16 93.34 71.57 87.84 87.25 31.21 72.15 63.04 73.32 78.14
MaPLe [18] 81.79 (75.58) 72.33 96.00 92.83 83.57 97.00 85.33 48.40 75.53 71.33 92.33 85.03
PSRC [19] 82.87 (77.90) 73.17 96.07 93.67 83.83 97.60 87.50 50.83 77.23 72.73 92.43 86.47
LLaMP 83.81 (78.50) 73.49 97.08 94.21 86.07 98.06 87.62 56.07 77.02 74.17 91.31 86.84
Table 2: Few shot classification results with 16 shots. Numbers in the bracket indicate the average performance over 1/2/4/8/16 shots.
@@ 4.3 Ablation Study
Noun phrases are parsed from the LLM’s responses the prompt of “Describe [STH]”. We then use the template, “A photo of [STH] with [NP]” to generate the NP-augmented text embedding for CLIP. We take the average of all augmented embeddings for classification. In Tab. 3 we show that even ordinary CLIP can benefit from incorporating LLMs’ knowledge.
@@
Refer to caption four engines, turboprop aircraft, large vertical fin Refer to caption
@@
Refer to caption a cluster, rectangular structures, flat roofs, straight lines Refer to caption
@@
Visualizations. In Fig. 4, we visualize the gradient heatmap of input images from FGVCAircraft and EuroSAT, through GradCAM [35]. The figure shows that, LLaMP can capture distinctive features that matches LLM’s description.
@@ 5 Conclusion
Acknowledgment
This research was supported, in part, by the Office of Naval Research under grant #N00014-21-1-2802.
Supplementary Material
Introduction
In the supplementary material, we provide extra discussions that did not fit in the main paper due to the space limitation, including i) Ablation study on the textual augmentation described in Sec 3.3; ii) Few-shot classification results on 8/4/2/1 shots with comparisons with previous methods.
@@ Introduction
In addition to the 16-shot classification result reported in the main paper, we present few-shot classification results with with 8/4/2/1 numbers of shots in Tab. 9 and compare LLaMP against previous baseline models.